# SparseSwaps: Tractable LLM Pruning Mask Refinement at Scale

### v1

@@ 2.2 The SparseSwaps algorithm
Building upon the preceding observations, we present our complete algorithm. The method takes as input a weight matrix $W\in\mathbb{R}^{d_{out}\times d_{in}}$, the Gram matrix $G=XX^{\top}\in\mathbb{R}^{d_{in}\times d_{in}}$ (accumulated during calibration), and a warmstart pruning mask $M^{\text{init}}\in\{0,1\}^{d_{out}\times d_{in}}$ that already satisfies the desired sparsity constraints, e.g., obtained from Wanda (Sun et al., 2024) or RIA (Zhang et al., 2024).
@@
2:Improved pruning mask $M$
@@
4:for $i=1$ to $d_{out}$ do $\triangleright$ Process each row independently
5:  $w\leftarrow W_{i,:}$, $m\leftarrow M_{i,:}$ $\triangleright$ Extract row weights and mask
6:  $\mathcal{P}\leftarrow\{j:m_{j}=0\}$, $\mathcal{U}\leftarrow\{j:m_{j}=1\}$ $\triangleright$ Pruned and unpruned sets
7:  $c\leftarrow G\cdot((\mathbbm{1}-m)\odot w)$ $\triangleright$ Initialize correlation vector
8:  for $t=1$ to $T_{\max}$ do
9:   $(p^{*},u^{*})\leftarrow\arg\min_{(p,u)}\;\Delta L_{u,p}$ $\triangleright$ Best swap via Equation 3
10:   if $\Delta L_{u^{*},p^{*}}<0$ then $\triangleright$ Swap improves objective
11:     $m_{p^{*}}\leftarrow 1$, $m_{u^{*}}\leftarrow 0$ $\triangleright$ Perform swap
12:     $\mathcal{P}\leftarrow(\mathcal{P}\setminus\{p^{*}\})\cup\{u^{*}\}$, $\mathcal{U}\leftarrow(\mathcal{U}\setminus\{u^{*}\})\cup\{p^{*}\}$
13:     $c\leftarrow c+w_{u^{*}}G_{:,u^{*}}-w_{p^{*}}G_{:,p^{*}}$ $\triangleright$ Update correlation vector
14:   else
15:     break $\triangleright$ Local optimum reached
16:   end if
@@
1-Swap optimization (Lines 6-15): We iteratively find the swap $(p^{*},u^{*})$ minimizing $\Delta L_{u,p}$ (cf. Equation 3) among feasible pairs, evaluating each candidate in $\mathcal{O}(1)$ time. If $\Delta L_{u^{*},p^{*}}<0$, we accept the swap and update the correlation vector via Equation 4; otherwise we terminate. At all times, the swaps are appropriately constrained: per-row sparsity allows any swap maintaining $|\mathcal{P}|$ constant, while $N\text{:}M$ sparsity restricts swaps to within the same $N\text{:}M$ blocks.
@@
In practice, several factors further reduce runtime. First, we find that even $T_{\max}=1$ or $T_{\max}=2$ can drastically reduce the local pruning error; values around $T_{\max}=25$ often suffice to significantly lower model perplexity, with diminishing returns beyond $T_{\max}=100$. Second, row-wise processing can be batched and vectorized, enabling parallel swap cost computations and mask updates, and rows can be distributed across GPUs if needed. Third, the Gram matrix $G$ is computed once per layer and shared across all rows, and several summands of Equation 3 can be similarly precomputed once per layer.
@@ 3 Experimental Results
We outline our general experimental approach, detailing datasets, architectures, and metrics. Our code is publicly available at github.com/ZIB-IOL/SparseSwaps. Our study focuses on language modeling within Natural Language Processing (NLP). We use pretrained models from HuggingFace (Wolf et al., 2020), specifically LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-2-9B (Riviere et al., 2024), Yi-1.5-9B (Young et al., 2025), DeepSeek-7B-base (Bi et al., 2024), and Qwen2.5-7B (Yang et al., 2025). For calibration, we randomly draw sequences of 2048 tokens from the C4 dataset (Raffel et al., 2020). For validation, we similarly pick 100 sequences from the validation split. The model performance is assessed via perplexity on the WikiText dataset (Merity et al., 2016) and zero-shot accuracy on the EleutherAI evaluation set (Gao et al., 2023). Following Sun et al. (2023), we prune all linear layers, excluding the embedding and final linear head, with uniform sparsity allocation across layers. We provide experiments for unstructured and semi-structured sparsity patterns (Mishra et al., 2021). We use multiple random seeds throughout our experiments.
@@ 3.1 Mask refinement at scale
SparseSwaps consistently improves state-of-the-art methods. Table 2 summarizes the main results and reports perplexity (upper half, lower is better) and zero-shot accuracy (lower half, higher is better) for warmstart masks (Wanda, RIA) as well as their refinements using DSnoT and SparseSwaps. For both 60% unstructured and 2:4 semi-structured sparsity, SparseSwaps (with 100 1-swap iterations) consistently reduces perplexity and improves zero-shot accuracy over Wanda and RIA warm start masks. While DSnoT similarly yields improvements, it falls short of SparseSwaps. Note that we left the pruning criterion of DSnoT, which partially uses the Wanda saliency, unchanged, even when using RIA warmstart. For unstructured RIA, we report results when enforcing a per-row sparsity constraint; while RIA yields good (and slightly better) results when enforcing truely unstructured sparsity, we decided to include the results for the per-row setting as this allows direct refinement of the mask with SparseSwaps and DSnoT.
SparseSwaps successfully optimizes the per-layer pruning loss. Figure 1 shows the per-layer reductions in local pruning error relative to a Wanda Warmstart, grouping layers by their corresponding Transformer block of LLaMA-3.1-8B. We observe drastic improvements of close to 70% compared to Wanda, demonstrating that SparseSwaps is able to successfully optimize the local loss. The attn.o_proj seems to consistently benefit the most across blocks, with reductions of the objective in Equation 1 ranging between 40%-60%.
Large local error reductions do not always imply reduced perplexity. From Table 2 we observe substantial perplexity gains, especially when sparsity more strongly degrades model quality (cf. Table 4 in the appendix, which shows more drastic improvements when using magnitude pruning, which more strongly degrades model quality). In contrast, when quality is less affected (e.g., at 50% sparsity where Wanda performs well), SparseSwaps yields limited perplexity gains despite significant local error reductions: Table 1 reports perplexity and average relative error reduction (%) versus the number of 1-swap iterations. Zero iterations correspond to the Wanda warm start; one or more iterations correspond to SparseSwaps from Wanda. At 50% sparsity, a single 1-swap iteration lowers relative error by 6.34%, and 200 iterations by nearly 40%, yet perplexity does not improve, but rather slightly increases. This suggests further reducing local error can overfit the calibration data and may not translate to better perplexity, although we note that the perplexity increase is relatively small. These results emphasize that while the reduction of local error is a useful proxy for perplexity reduction when pruning has a higher negative impact on the model, the local error of Equation 1 remains an approximation to the reconstruction error of the entire model.
Table 1: LLaMA-3.1-8B: Perplexity ($\downarrow$) and mean relative reduction in pruning error ($\uparrow$) versus number of 1-swap iterations for 50% and 60% unstructured sparsity using Wanda warmstart.
Number of 1-swap iterations
Sparsity Metric 0 1 2 5 10 25 50 100 200
50% Avg. rel. error reduction (%) 0.00 6.34 8.77 12.51 16.38 23.52 30.04 36.48 38.95
Perplexity 10.13 10.31 10.40 10.41 10.39 10.38 10.27 10.30 10.34
60% Avg. rel. error reduction (%) 0.00 8.04 11.04 15.34 19.64 26.92 33.58 39.99 43.74
Perplexity 21.52 21.26 21.51 21.17 21.01 20.38 19.74 18.96 19.17
@@
Perplexity $\downarrow$ LLaMA-3.1 Gemma-2 Yi-1.5 DeepSeek Qwen2.5
Method Sparsity 8B 9B 9B 7B 7B
Wanda 60% 21.94 16.74 11.40 11.41 13.75
+ DSnoT 60% 21.94 16.69 11.38 11.40 13.75
+ SparseSwaps 60% 19.75 16.01 10.07 10.93 13.16
RIA 60% 19.73 16.19 10.73 11.80 12.63
+ DSnoT 60% 19.73 16.22 10.73 11.80 12.63
+ SparseSwaps 60% 18.47 15.44 9.98 10.79 12.47
Wanda 2:4 24.82 17.45 11.76 11.77 14.53
+ DSnoT 2:4 22.79 16.79 10.84 11.70 14.40
+ SparseSwaps 2:4 20.17 16.30 10.73 11.70 13.95
RIA 2:4 23.96 16.88 11.29 12.03 13.58
+ DSnoT 2:4 24.26 16.82 10.57 12.03 13.85
+ SparseSwaps 2:4 20.90 16.33 10.50 11.80 13.28
Accuracy $\uparrow$ LLaMA-3.1 Gemma-2 Yi-1.5 DeepSeek Qwen2.5
Method Sparsity 8B 9B 9B 7B 7B
Wanda 60% 48.18% 63.39% 53.59% 50.74% 59.26%
+ DSnoT 60% 48.18% 63.49% 53.79% 50.75% 59.26%
+ SparseSwaps 60% 50.78% 63.84% 54.84% 51.02% 60.15%
RIA 60% 49.56% 64.37% 52.81% 50.92% 59.84%
+ DSnoT 60% 49.56% 64.43% 52.96% 50.83% 59.81%
+ SparseSwaps 60% 51.02% 64.32% 54.45% 51.47% 61.22%
Wanda 2:4 46.80% 63.73% 52.58% 51.02% 59.52%
+ DSnoT 2:4 47.01% 63.66% 52.16% 50.78% 59.09%
+ SparseSwaps 2:4 48.83% 64.70% 52.43% 50.36% 59.92%
RIA 2:4 47.87% 63.87% 52.68% 51.22% 58.66%
+ DSnoT 2:4 47.13% 64.17% 51.36% 49.86% 59.72%
+ SparseSwaps 2:4 49.90% 64.60% 52.30% 51.46% 60.31%
Figure 1: Per-layer relative reduction in local pruning error compared to Wanda. The plot shows result for LLaMA-3.1-8B, 60% unstructured sparsity and 100 1-swap iterations.
3.2 Efficiency and hyperparameter ablations
Resource requirements. SparseSwaps is more resource-intensive than DSnoT and, as a drop-in refinement, requires at least the resources of the chosen warm-start method. Beyond that, SparseSwaps needs memory to store the Gram matrix $G\in\mathbb{R}^{d_{in}\times d_{in}}$ (once per layer) and the correlation vector $c\in\mathbb{R}^{d_{in}}$ (per row), and compute to perform the 1-swaps; see the preceding section for the theoretical complexity. While we have argued in the introduction that the additional compute can be justified when amortized over many LLM inference requests, we note that the overhead grows only linearly with the number of 1-swap iterations $T_{\max}$. Table 1 shows that few iterations already yield substantial gains in both perplexity and local error reduction, especially at higher sparsity.
Table 3 reports wall-clock times for pruning LLaMA-3.1-8B to 60% sparsity on a single H100 GPU. The $T_{\max}=0$ baseline includes calibration data sampling, Wanda pruning, Gram matrix computation, and evaluation; each additional iteration of SparseSwaps adds a relatively small overhead. For comparison, Wanda and SparseGPT take approximately 4 and 10 minutes, respectively. We note that our implementation can be further optimized and that the algorithm is fully parallelizable across rows.
Table 3: Wall-clock time for applying SparseSwaps to LLaMA-3.1-8B at 60% sparsity on a single H100 GPU.
$T_{\max}$ 0 1 2 5 10 25
Wall-clock time 8m15s 10m17s 12m7s 17m20s 26m13s 52m29s
Effect of the number of reconstruction samples. Figure 2 in the appendix shows the perplexity versus the number of reconstruction samples for 50% and 60% unstructured sparsity when using Wanda as well as SparseSwaps with a Wanda warmstart. We observe that the perplexity decreases drastically when using more samples, which leads to SparseSwaps slightly outperforming Wanda for 50% sparsity, despite its advantage typically being larger at higher sparsity. We emphasize that the number of reconstruction samples does not affect SparseSwaps’s swap evaluation efficiency: the Gram matrix $G=XX^{\top}$ has fixed size $d_{in}\times d_{in}$ regardless of $B$.
4 Conclusion
We revisited the mask selection problem for post-training pruning and showed that it can be made substantially more tractable, even at LLM scale. We observed that row decoupling via equal per-row sparsity yields independent subproblems, and that individual 1-swaps can be evaluated in $\mathcal{O}(1)$ time using the Gram matrix $G=XX^{\top}$. This enables tractable optimization of the true row-wise quadratic loss on GPUs. The resulting method, SparseSwaps, is warm-start agnostic, nearly hyperparameter-free, and scalable. It consistently reduces per-layer pruning error and improves perplexity and zero-shot accuracy across modern GPT architectures.
Our work is not without limitations. While per-row sparsity is not necessarily detrimental for LLMs, our approach is restricted to that setting and only partially adapts to truly unstructured sparsity; in its current form, the algorithm can handle unstructured sparsity but cannot reallocate sparsity levels across rows. Furthermore, runtime and memory remain non-trivial for large architectures.
Acknowledgments
This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG) as well as by the German Federal Ministry of Research, Technology and Space (fund number 16IS23025B).
Appendix A Appendix
A.1 Further Results

### v2

@@ 2.2 The SparseSwaps Algorithm
Building upon the preceding observations, we present our complete algorithm. The method takes as input a weight matrix $W\in\mathbb{R}^{d_{out}\times d_{in}}$, the Gram matrix $G=XX^{\top}\in\mathbb{R}^{d_{in}\times d_{in}}$ (accumulated during calibration), and a warmstart pruning mask $M^{\text{init}}\in\{0,1\}^{d_{out}\times d_{in}}$ that already satisfies the desired sparsity constraints, e.g., obtained from Wanda (Sun et al., 2024) or RIA (Zhang et al., 2024a).
@@
0: Improved pruning mask $M$
@@
2:for $i=1$ to $d_{out}$ do
3:  $w\leftarrow W_{i,:}$, $m\leftarrow M_{i,:}$
4:  $\mathcal{P}\leftarrow\{j:m_{j}=0\}$, $\mathcal{U}\leftarrow\{j:m_{j}=1\}$
5:  $c\leftarrow G\cdot((\mathbbm{1}-m)\odot w)$
6:  for $t=1$ to $T_{\max}$ do
7:   $(p^{*},u^{*})\leftarrow\arg\min_{(p,u)}\;\Delta L_{u,p}$ Best swap
8:   if $\Delta L_{u^{*},p^{*}}<0$ then
9:    $m_{p^{*}}\leftarrow 1$, $m_{u^{*}}\leftarrow 0$ Perform swap
10:    $\mathcal{P}\leftarrow(\mathcal{P}\setminus\{p^{*}\})\cup\{u^{*}\}$, $\mathcal{U}\leftarrow(\mathcal{U}\setminus\{u^{*}\})\cup\{p^{*}\}$
11:    $c\leftarrow c+w_{u^{*}}G_{:,u^{*}}-w_{p^{*}}G_{:,p^{*}}$
12:   else
13:    break Local optimum reached
14:   end if
15:  end for
16:  $M_{i,:}\leftarrow m$ Store optimized row
@@
1-Swap optimization (Lines 6-15): We iteratively find the swap $(p^{*},u^{*})$ minimizing $\Delta L_{u,p}$ (cf. Equation 5) among feasible pairs, evaluating each candidate using only lookups. If $\Delta L_{u^{*},p^{*}}<0$, we accept the swap and update the correlation vector via Equation 6; otherwise we terminate. At all times, the swaps are appropriately constrained: per-row sparsity allows any swap maintaining $|\mathcal{P}|$ constant, while $N\text{:}M$ sparsity restricts swaps to within the same $N\text{:}M$ blocks.
@@
Convergence.
Intuitively, since each accepted swap strictly decreases the per-row loss $L$ and $L\geq 0$, the algorithm must terminate at a 1-swap local optimum, i.e., a mask from which no single swap can reduce the loss. We state this simple observation below and defer the formal statement and proof to the appendix.
Proposition 2.1 (Convergence, informal).
Suppose Algorithm 1 terminates when no swap improves the loss by more than $\varepsilon\geq 0$. Then, starting from initial per-row loss $L_{0}$, it performs at most $\mathcal{O}(L_{0}/\varepsilon)$ swaps to reach a $\varepsilon$-1-swap local optimum.
In practice, several factors further reduce the runtime. First, we find that even $T_{\max}=1$ or $T_{\max}=2$ can drastically reduce the local pruning error; values around $T_{\max}=25$ often suffice to significantly lower model perplexity, with diminishing returns beyond $T_{\max}=100$. Second, row-wise processing can be batched and vectorized, enabling parallel swap cost computations and mask updates, and rows can be distributed across GPUs if needed. Third, the Gram matrix $G$ is computed once per layer and shared across all rows, and several summands of Equation 5 can be similarly precomputed once per layer.
@@ 3 Experimental Results
We outline our general experimental approach, detailing datasets, architectures, and metrics. Our code is publicly available at github.com/ZIB-IOL/SparseSwaps. We focus on language modeling within Natural Language Processing (NLP). We use pretrained models from HuggingFace (Wolf et al., 2020), specifically LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-2-9B (Riviere et al., 2024), Yi-1.5-9B (Young et al., 2025), DeepSeek-7B-base (Bi et al., 2024), and Qwen2.5-7B (Yang et al., 2025). For calibration, we randomly draw sequences of 2048 tokens from the C4 dataset (Raffel et al., 2020). For validation, we similarly pick 100 sequences from the validation split. The model performance is assessed via perplexity on the WikiText dataset (Merity et al., 2016) and zero-shot accuracy on the EleutherAI evaluation set (Gao et al., 2023). Following Sun et al. (2024), we prune all linear layers, excluding the embedding and final linear head, with uniform sparsity allocation across layers. We provide experiments for unstructured and semi-structured sparsity patterns (Mishra et al., 2021). We use multiple random seeds throughout our experiments.
@@ 3.1 Mask Refinement at Scale
Perplexity $\downarrow$ Accuracy $\uparrow$
Method Sparsity LLaMA-3.1 Gemma-2 Yi-1.5 DeepSeek Qwen2.5 LLaMA-3.1 Gemma-2 Yi-1.5 DeepSeek Qwen2.5
8B 9B 9B 7B 7B 8B 9B 9B 7B 7B
Wanda 60% 21.94 16.74 11.40 11.41 13.75 48.18% 63.39% 53.59% 50.74% 59.26%
+ DSnoT 60% 21.94 16.69 11.38 11.40 13.75 48.18% 63.49% 53.79% 50.75% 59.26%
+ SparseSwaps 60% 19.75 16.01 10.07 10.93 13.16 50.78% 63.84% 54.84% 51.02% 60.15%
RIA 60% 19.73 16.19 10.73 11.80 12.63 49.56% 64.37% 52.81% 50.92% 59.84%
+ DSnoT 60% 19.73 16.22 10.73 11.80 12.63 49.56% 64.43% 52.96% 50.83% 59.81%
+ SparseSwaps 60% 18.47 15.44 9.98 10.79 12.47 51.02% 64.32% 54.45% 51.47% 61.22%
Wanda 2:4 24.82 17.45 11.76 11.77 14.53 46.80% 63.73% 52.58% 51.02% 59.52%
+ DSnoT 2:4 22.79 16.79 10.84 11.70 14.40 47.01% 63.66% 52.16% 50.78% 59.09%
+ SparseSwaps 2:4 20.17 16.30 10.73 11.70 13.95 48.83% 64.70% 52.43% 50.36% 59.92%
RIA 2:4 23.96 16.88 11.29 12.03 13.58 47.87% 63.87% 52.68% 51.22% 58.66%
+ DSnoT 2:4 24.26 16.82 10.57 12.03 13.85 47.13% 64.17% 51.36% 49.86% 59.72%
+ SparseSwaps 2:4 20.90 16.33 10.50 11.80 13.28 49.90% 64.60% 52.30% 51.46% 60.31%
SparseSwaps consistently improves state-of-the-art methods. Table 1 summarizes the main results and reports perplexity (upper half, lower is better) and zero-shot accuracy (lower half, higher is better) for warmstart masks (Wanda, RIA) as well as their refinements using DSnoT and SparseSwaps. For both 60% unstructured and 2:4 semi-structured sparsity, SparseSwaps (with 100 1-swap iterations) consistently reduces perplexity and improves zero-shot accuracy over Wanda and RIA warmstart masks. While DSnoT similarly yields improvements, it falls short of SparseSwaps. Note that we left the pruning criterion of DSnoT, which partially uses the Wanda saliency, unchanged, even when using RIA warmstart. For unstructured RIA, we report results when enforcing a per-row sparsity constraint; while RIA yields good (and slightly better) results when enforcing truly unstructured sparsity, we decided to include the results for the per-row setting as this allows direct refinement of the mask with SparseSwaps and DSnoT.
SparseSwaps successfully optimizes the per-layer pruning loss. Figure 1 shows the per-layer reductions in local pruning error relative to a Wanda warmstart, grouping layers by their corresponding Transformer block of LLaMA-3.1-8B. We observe drastic improvements of close to 70% compared to Wanda, demonstrating that SparseSwaps is able to successfully optimize the local loss. The attn.o_proj seems to consistently benefit the most across blocks, with reductions of the objective in Equation 1 ranging between 40% and 60%.
@@
Large local error reductions do not always imply reduced perplexity. From Table 1 we observe substantial perplexity gains. Surprisingly, when the model quality is less affected by pruning (e.g., at 50% sparsity where Wanda performs well), SparseSwaps yields limited perplexity gains despite significant local error reductions: Table 3 reports perplexity and average relative error reduction (%) versus the number of 1-swap iterations. Zero iterations correspond to the Wanda warmstart; one or more iterations correspond to SparseSwaps from Wanda. At 50% sparsity, a single 1-swap iteration lowers relative error by 6.34%, and 100 iterations by close to 40%, yet perplexity does not improve, but rather slightly increases. This suggests further reducing local error can overfit the calibration data and may not translate to better perplexity, although we note that the perplexity increase is relatively small. These results emphasize that while the reduction of local error is a useful proxy for perplexity reduction when pruning has a higher negative impact on the model, the local error of Equation 1 remains only an approximation to the reconstruction error of the entire model. Table 2 shows more drastic improvements when using magnitude pruning, further highlighting that the impact of SparseSwaps is most pronounced when model degradation is high.
Table 3: LLaMA-3.1-8B: Mean relative reduction in pruning error ($\uparrow$) and perplexity ($\downarrow$) versus number of 1-swap iterations for 50% and 60% unstructured sparsity using Wanda warmstart.
Number of 1-swap iterations
Sparsity Metric 0 1 2 5 10 25 50 100
50% Error reduction (%) 0.00 6.34 8.77 12.51 16.38 23.52 30.04 36.48
Perplexity 10.13 10.31 10.40 10.41 10.39 10.38 10.27 10.30
60% Error reduction (%) 0.00 8.04 11.04 15.34 19.64 26.92 33.58 39.99
Perplexity 21.52 21.26 21.51 21.17 21.01 20.38 19.74 18.96
SparseSwaps is robust to warmstart quality. Table 4 shows that weaker warmstarts yield larger error reductions (58–63% from magnitude vs. 37–43% from Wanda), demonstrating that SparseSwaps is not sensitive to warmstart quality and can recover from poor initializations.
Table 4: Average relative reduction in local pruning error (%) for different warmstart methods at 60% sparsity. Weaker warmstarts yield larger reductions.
Warmstart LLaMA-3.1-8B Gemma-2-9B DeepSeek-7B
Magnitude 62.69% 58.13% 58.10%
Wanda 43.29% 40.45% 36.96%
3.2 Efficiency and Hyperparameter Ablations
Resource requirements. SparseSwaps is more resource-intensive than DSnoT and, as a drop-in refinement, requires at least the resources of the chosen warmstart method. Beyond that, SparseSwaps needs memory to store the Gram matrix $G\in\mathbb{R}^{d_{in}\times d_{in}}$ (once per layer) and the correlation vector $c\in\mathbb{R}^{d_{in}}$ (per row), as well as the compute to perform the 1-swaps; see the preceding section for the theoretical complexity. While we have argued in the introduction that the additional compute can be justified when amortized over many LLM inference requests, we note that the overhead grows only linearly with the number of 1-swap iterations $T_{\max}$. Table 3 shows that few iterations already yield substantial gains in both perplexity and local error reduction, especially at higher sparsity.
Table 5: Wall-clock time in minutes for applying SparseSwaps to LLaMA-3.1-8B at 60% sparsity on a single H100 GPU. The $T_{\max}=0$ baseline includes calibration data sampling, Wanda pruning, Gram matrix computation, and evaluation.
$\mathbf{T_{\max}}$ 0 1 2 5 10 25
Wall-clock time 8m 10m 12m 17m 26m 52m
Table 5 further reports wall-clock times for pruning LLaMA-3.1-8B to 60% sparsity on a single H100 GPU. The $T_{\max}=0$ baseline includes calibration data sampling, Wanda pruning, Gram matrix computation, and evaluation; each additional iteration of SparseSwaps adds a relatively small overhead. For comparison, Wanda and SparseGPT take approximately 4 and 10 minutes, respectively. We note that our implementation can be further optimized and that the algorithm is fully parallelizable across rows.
Effect of the number of reconstruction samples. Figure 2 in the appendix shows the perplexity versus the number of reconstruction samples for 50% and 60% unstructured sparsity when using Wanda as well as SparseSwaps with a Wanda warmstart. We observe that the perplexity decreases drastically when using more samples, which leads to SparseSwaps slightly outperforming Wanda for 50% sparsity, despite its advantage typically being larger at higher sparsity. We emphasize that the number of reconstruction samples does not affect SparseSwaps’ swap evaluation efficiency: the Gram matrix $G=XX^{\top}$ has fixed size $d_{in}\times d_{in}$ regardless of $B$.
4 Conclusion
We revisited the mask selection problem for post-training pruning and showed that it can be made substantially more tractable, even at LLM scale. We observed that row decoupling via equal per-row sparsity yields independent subproblems, and that individual 1-swaps can be evaluated in $\mathcal{O}(1)$ time using the Gram matrix $G=XX^{\top}$. This enables tractable optimization of the true row-wise quadratic loss on GPUs. The resulting method, SparseSwaps, is warmstart-agnostic, nearly hyperparameter-free, and scalable. It consistently reduces per-layer pruning error and improves perplexity and zero-shot accuracy across modern GPT architectures.
Our work is not without limitations. While per-row sparsity is not necessarily detrimental for LLMs, our approach is restricted to that setting and only partially adapts to truly unstructured sparsity; in its current form, the algorithm can handle unstructured sparsity but cannot reallocate sparsity levels across rows. A reallocation of sparsity between individual rows might pose an interesting direction for further research. Further, the focus of our work is to demonstrate that exact mask refinement is realizable at LLM scale and can improve existing mask-selection methods. Our objective is to refine pruning masks post-hoc. There exists a variety of methods that optimize both the mask as well as the remaining, non-pruned weights jointly. In contrast, SparseSwaps focuses on mask optimization and is complementary to such approaches, which can in principle be combined with SparseSwaps to refine intermediate pruning steps. We leave such extensions for future work.
Acknowledgments
This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/2, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG) as well as by the German Federal Ministry of Research, Technology and Space (research campus Modal, fund number 05M14ZAM, 05M20ZBM) and the VDI/VDE Innovation + Technik GmbH (fund number 16IS23025B).
Appendix A Appendix
A different (but in practice slightly less efficient) perspective on the reduction of the computational complexity through the Gram matrix is through the unitary invariance of the Frobenius norm used in our pruning objective: for any matrix $A$ and unitary matrix $U$ (i.e., $U^{-1}=U^{\top}$), we have ${\left\|AU\right\|}_{F}={\left\|A\right\|}_{F}$. This property enables significant computational savings through Singular Value Decomposition (SVD) compression. Precisely, let $X=U\Sigma V^{\top}$ be the SVD of calibration data $X\in\mathbb{R}^{d_{in}\times B}$. Since $B>d_{in}$, we can write $\Sigma=[\Sigma^{\prime}\mid 0]$ with $\Sigma^{\prime}\in\mathbb{R}^{d_{in}\times d_{in}}$ containing the singular values on its diagonal. The compressed representation is simply $X^{\prime}=U\Sigma^{\prime}\in\mathbb{R}^{d_{in}\times d_{in}}$. Letting $w_{p}=w-m\odot w$ for brevity, the key insight is that pruning decisions remain equivalent under this compression:
${\left\|w_{p}X\right\|}_{F}^{2}$ $={\left\|w_{p}U\Sigma V^{\top}\right\|}_{F}^{2}={\left\|w_{p}U\Sigma\right\|}_{F}^{2}$
$={\left\|w_{p}U[\Sigma^{\prime}\mid 0]\right\|}_{F}^{2}={\left\|w_{p}U\Sigma^{\prime}\right\|}_{F}^{2}={\left\|w_{p}X^{\prime}\right\|}_{F}^{2},$
A.1 Convergence to a Local Optimum
We formalize the convergence guarantee for SparseSwaps. Recall that for a single row with pruned index set $\mathcal{P}$ and unpruned set $\mathcal{U}$, the per-row loss is $L=\|r\|_{2}^{2}$ where $r=\sum_{j\in\mathcal{P}}w_{j}\phi_{j}$. A 1-swap exchanges one index between $\mathcal{P}$ and $\mathcal{U}$, preserving the sparsity level.
Definition A.1 ($\varepsilon$-1-swap local optimum).
A mask $m$ (equivalently, its pruned set $\mathcal{P}$) is an $\varepsilon$-1-swap local optimum if
$\min_{u\in\mathcal{U},p\in\mathcal{P}}\Delta L_{u,p}\geq-\varepsilon,$
i.e., no single swap can reduce the per-row loss $L$ by more than $\varepsilon$.
Proposition A.2 (Convergence).
Let $m_{0}$ be the initial mask with per-row loss $L_{0}$. Suppose SparseSwaps terminates when the best swap satisfies $\Delta L_{u^{*},p^{*}}\geq-\varepsilon$ for some $\varepsilon\geq 0$. Then SparseSwaps performs at most $T\leq\lceil L_{0}/\varepsilon\rceil$ swap operations before reaching an $\varepsilon$-1-swap local optimum.
Proof.
At any iteration $t$ that does not terminate, the algorithm accepts a swap with $\Delta L_{u^{*},p^{*}}<-\varepsilon$ by definition of the stopping rule. Since $L_{t+1}=L_{t}+\Delta L_{u^{*},p^{*}}$, this yields the strict decrease $L_{t+1}<L_{t}-\varepsilon$. Unrolling over $T$ accepted swaps gives $L_{T}<L_{0}-T\varepsilon$. Since $L$ is a squared Euclidean norm, we have $L_{T}\geq 0$, which implies $T<L_{0}/\varepsilon$, and hence $T\leq\lceil L_{0}/\varepsilon\rceil$. At termination, we have $\min_{(u,p)}\Delta L_{u,p}\geq-\varepsilon$, so the final mask is an $\varepsilon$-1-swap local optimum. ∎
A.2 Further Results
Figure 2 shows the perplexity versus the number of reconstruction samples for 50% and 60% unstructured sparsity when using Wanda as well as SparseSwaps with a Wanda warmstart. We further ran an ablation study comparing Wanda and SparseSwaps when varying the sequence length while keeping the number of calibration samples fixed at 128. Both methods were equally impacted by the sequence length, with the overall impact being rather minor in terms of final perplexity.